Exercise 2-2 Apply Overhead Cost to Jobs [LO2-2]

Luthan Company uses a plantwide predetermined overhead rate of $22.30 per direct labor-hour. This predetermined rate was based on a cost formula that estimated $267,600 of total manufacturing overhead cost for an estimated activity level of 12,000 direct labor-hours.

The company incurred actual total manufacturing overhead cost of $267,000 and 11,500 total direct labor-hours during the period.

Required:

Determine the amount of manufacturing overhead cost that would have been applied to all jobs during the period.
Answer #1

Answer : the amount of manufacturing overhead cost that would have been applied to all jobs during the period is $256,450

Explanation :

Step 1:Pre determined overhead rate per direct labor hour is already given = $22.30

It is calculated as

(Pre determined overhead rate = estimated manufacturing overhead cost /estimated activity level

=$267,600/12000=22.30)

Step 2: multiply the actual labor hours incurred during the period with the pre determined overhead rate to get the manufacturing overhead cost  applied to all the jobs .

= 11500*22.30=$256,450 (answer)